An account invested in a money market fund grew from ​$73,334.75 to ​$73,630.70 in a month. What was the interest rate to the nearest​ tenth? What was the interest​ rate?
Answer #1

Answer:

Initial Amount P=$73334.75

Final Amount F=$73630.70

So Interest Rate per Month =(F-P)/P=(73630.70-73334.75)/73334.75=0.404% per month

Nominal Annual Interest rate =12*0.404%=4.84%

Effective annual interest rate=(1+o.404%)^12-1=4.95%